sdbfsd:

A substance decays so that the amount A of the substance left after t years is given by: A = A0 · (0.9)t, where A0 is the original amount of the substance. What is the half-life (the amount of time that it takes to decay to half the original amount) of this substance rounded to the nearest tenth of a yea

sdbfsd:

help

sdbfsd:

6.6 years 7.5 years 6.1 years 5.6 years

sdbfsd:

those are the answers

Vocaloid:

A = A0 · (0.9)^t At a substance's half life, the final amount A equals (1/2)A0 so substituting (1/2)A0 for A on the left side of the equation: (1/2)A0 = A0 · (0.9)^t cancel out A0 from both sides, then solve for t